Circular No. 8369

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

NOVA SCORPII 2004 G. Pojmanski, Warsaw University Astronomical Observatory, reports

the discovery of a variable star on ‘All Sky Automated Survey’ images obtained with a 65-mm-aperture (180-mm-focal-length f/2.8) telephoto lens at Las Campanas on July 3.146 (at V = 11.98) and 5.239 UT (V = 11.19), with the position of the variable given as α = 17^{h}12^{m}51^{s}.3, δ = −30^{o}56′39′′ (equinox 2000.0; CCD pixel size 14′′.8). Pojmanski reports that nothing was visible at this location on an ASAS image taken on June 27.219 (limiting mag V ≈ 14.0). H. Yamaoka, Kyushu University, reports his measures for the new object from an image taken by G. Bolt (Craigie, Australia), yielding position end figures 51^{s}.28, 37′′.6 (82 UCAC-2 reference stars; fitting error ∼ 0′′.15 for each coordinate); Yamaoka adds that the nearest star visible in DSS images (and listed in the GSC 2.2 and 2MASS catalogues) is ∼ 1′′ from this position. D. Nogami, Hida Observatory, Kyoto University, finds that there is no known x-ray source within 30′′ of the ASAS position, which suggests that it is not a dwarf nova. Nogami adds that a spectrum of the new object, taken by M. Fujii (Bisei, Okayama, Japan) on July 6.51, shows prominent Balmer emission lines (FWHM ∼ 500 km/s for H_{α}, with possible P-Cyg profile); Na D absorption on a flat continuum is also clearly detected, which suggests moderate reddening. Nogami suggests that the spectrum resembles that of the slow nova V723 Cas (= N Cas 1995) during the premaximum halt phase (Iijima et al. 1998, A.Ap. 338, 1006). Additional magnitude estimates (partially forwarded by Yamaoka; visual unless noted otherwise): July 4.985, [11.8 (W. Liller, Viña del Mar, Chile, Technical Pan film + orange filter); 5.645, 11.3 (Bolt, unfiltered CCD; 0.06-mag rise in 2.7 hr); 6.052, R = 11.0 (P. Cacella, Brasilia, Brazil; CCD); 6.234, V = 10.92 (Pojmanski); 6.281, 11.1 (M. Linnolt, Honolulu, HI); 6.413, 10.8 (J. Bedient, Honolulu, HI); 6.463, B = 11.62, V = 10.83, R_{c} = 10.41, I_{c} = 10.00 (S. Kiyota, Tsukuba, Japan); 6.567, 10.9 (Y. Maeda, Nagasaki, Japan); 6.616, B = 11.44, V = 10.83, R_{c} = 10.33, I_{c} = 9.91 (H. Maehara, Saitama, Japan).

UNUSUAL CATACLYSMIC VARIABLE IN HERCULES CCD magnitudes by B. E. Schaefer, Louisiana State University, via the

Cerro Tololo 1.0-m telescope, for the variable on IAUC 8363: July 1.198 UT, U = 15.03, B = 15.87, V = 15.68, R = 15.24, I = 14.37; 3.250, 15.46, 16.31, 16.03, 15.44, 14.48; July 5.183, 15.20, 15.96, 16.03, 15.54, 14.52.

2004 July 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green